# REMI: Reconstructing Episodic Memory During Intrinsic Path Planning

### v1

@@ 3 Recalling MEC Representations from Sensory Observations
We first tested that sensory observations at the goal location can trigger retrieval of the corresponding grid cell (GC) representation through auto-association of GCs and spatially modulated cells (SMCs). Auto-association is expected to occur when the input pattern is incomplete or degraded. To evaluate this, we trained nine models with identical configurations, varying only the masking ratio $r_{\text{mask}}$ from 0.1 to 0.9. The masking ratio $r_{\text{mask}}$ specifies the maximum fraction of direction, speed, GC, and SMC inputs, as well as their initial hidden states, that are randomly set to zero during training. This simulates degraded observations and encourages the network to learn robust recall through auto-association. Each model was trained on randomly sampled short trajectories using a fixed $r_{\text{mask}}$, with new random masks generated for every trajectory and varying across time and cells. Masks were applied to both the inputs and initial states of GCs, SMCs, speed cells, and direction cells (Suppl. 5.2).
@@
Figure 2a shows the L2 distance between decoded and ground-truth positions over time. The network was queried for 5 seconds (100 timesteps), and all models successfully recalled the goal location with high accuracy. HPCs were identified as neurons with mean firing rates above 0.01 Hz and spatial information content (SIC) above 20 (Suppl. 2.1). The number of HPCs increased with $r_{\text{mask}}$, and location decoding using HPCs was performed only for models with more than 10 HPCs. We observe a trade-off in which higher $r_{\text{mask}}$ leads to more HPCs and improved decoding accuracy but reduces the network’s ability to recall sensory observations (Figure 2a,b).
@@ 4 Planning with Recalled Representations
The recall experiment above demonstrates that auto-association of spatially modulated cells (SMCs) and grid cells (GCs) through hippocampal place cells (HPCs) could enable the recall of all cells’ (GC, HPC, and SMC) representations using sensory cues. Among the recalled patterns, we propose that animals primarily use GCs for planning, as direct planning with HPCs or SMCs cannot take shortcut paths through unvisited regions. Auto-association removes the need for direct planning with SMCs or HPCs; sensory cues are required only to determine initial and target GC representations. Once recalled, a sequence of speed and direction signals drives GC activity from the initial to the target pattern, generating the planned path. Sensory patterns can then be reconstructed from intermediate GC states through HPCs, further reducing reliance on SMCs and HPCs during planning. These speed and direction signals are context-free and generalizable across environments, allowing planning strategies learned in one room to transfer to others. Finally, planning within GC representations may enable local transition rules to generalize to long-range path planning.
@@ 4.1 Decoding Displacement from Grid Cells
Consider a 1D case with $\phi_{c}$ and $\phi_{t}$ as the phase variables of the current and target locations in some module. The phase difference is $\Delta\phi=\phi_{t}-\phi_{c}$, and since $\phi_{c},\phi_{t}\in[0,2\pi)$, we have $\Delta\phi\in(-2\pi,2\pi)$. However, for vector-based navigation, we instead need $\Delta\phi^{*}$ such that $\phi_{t}=\left[\phi_{c}+\Delta\phi^{*}\right]_{2\pi}$, where $[\cdot]_{2\pi}$ is an element-wise modulo operation so that $\phi_{t}$ is defined on $[0,2\pi)$. Simply using $\Delta\phi$ directly is not sufficient because multiple wrapped phase differences correspond to the same phase $\phi_{t}$, but different physical positions on the torus. Therefore, we restrict $\Delta\phi$ to be defined on $(-\pi,\pi)$ such that the planning mechanism always selects the shortest path on the torus that points to the target phase. The decoded displacement in physical space is then $\hat{d}\in[-\ell/2,\ell/2]$.
@@
Figure 3: (a) The population response of a single grid module forms an $n$-dimensional torus, where multiple phase differences can connect the current and target phases. (b) Unwrapping phases into physical space yields $2^{n}$ candidate displacements; only $\phi_{t}^{*}$ lies within the principal domain $(-\pi,\pi)^{n}$. (c) A Markovian process asymptotically identifies the most likely next phase that moves closer to the target. (d) Renormalizing after each update produces a smooth trajectory from start to target. (e) Illustration of this process in 2D space. (f) Learned connectivity matrix of the planning RNN using only grid cells. (g) Planned trajectories for targets reachable within $10$ and $20$ seconds. Blue and red crosses mark start and target locations; the reference line shows the full trajectory for visualization. The dots represented the decoded locations. (h) A planning network connected to the full HC-MEC, receiving input only from GC and controlling speed and direction, drives SMC responses to update alongside GC, tracing a trajectory closely aligned with the planned GC path.
@@ 4.2 Sequential Planning
By repeatedly applying this update, the influence of the current and target phase spreads through phase space, eventually settling on an intermediate phase that connects start and target (Figure 3c). If the animal selects the phase with the highest value after each update and renormalizes the vector, this process traces a smooth trajectory toward the target (Figure 3d). This approach can be generalized to 2D phase spaces (Figure 3e). In essence, we propose that the animal can decompose long-range planning into a sequence of local steps by encoding a transition probability over phases in a matrix. A readout mechanism can then map these phase transitions into corresponding speeds and directions and subsequently update GC activity toward the target. We also note that this iterative planning process resembles search-based methods in robotics, such as RRT-Connect [67].
@@ 4.3 Combining Decoded Displacement from Multiple Scales
Given the ongoing debate about whether grid cells contribute to long-range planning [68, 69], we focus on mid-range planning, where distances are of the same order of magnitude as the environment size. Suppose two locations in 1D space are separated by a ground truth displacement $d\in\mathbb{R}_{+}$, bounded by half the largest scale ($\ell_{m}/2$). We can always find an index $k$ where $\ell_{1}/2,\dots,\ell_{k}/2\leq d<\ell_{k+1}/2<\dots,\ell_{m}/2$. Given $k$, we call scales $\ell_{k+1},\dots,\ell_{m}$ decodable and scales $\ell_{1},\dots,\ell_{k}$ undercovered. For undercovered scales, phase differences $\Delta\phi$ are wrapped around the torus at least one period of $(-\pi,\pi)$ and may point to the wrong direction. We thus denote phase difference from undercovered scales as $Z_{i}$. If we predict displacement by simply averaging the decoded displacements from all grid scales, the predicted displacement is:
$\hat{d}=\frac{\ell_{0}}{2\pi m}\cdot\left(\sum_{i=1}^{k}s^{i}\cdot Z_{i}+\sum_{i=k+1}^{m}s^{i}\cdot\Delta\phi_{i}\right)$
In 1D, the remaining distance after taking the predicted displacement is $d_{\text{next}}=d_{\text{current}}-\hat{d}$. For the predicted displacement to always move the animal closer to the target, meaning $d_{\text{next}}<d_{\text{current}}$, it suffices that $m>k+\frac{1-s^{-k}}{s-1}$ (see Suppl. 1.1). This condition is trivially satisfied in 1D for $s=e$, as $\frac{1-s^{-k}}{s-1}<1$ requiring only $m>k$. In 2D, where the optimal scaling factor $s=\sqrt{e}$, the condition tightens slightly to $m>k+1$. Importantly, as the animal moves closer to the target, more scales become decodable, enabling increasingly accurate predictions that eventually lead to the target. In 2D, planning can be decomposed into two independent processes along non-collinear directions. Although prediction errors in 2D may lead to a suboptimal path, this deviation can be reduced by increasing the number of scales or taking smaller steps along the decoded direction, allowing the animal to gradually approach the target with higher prediction accuracy.
@@ 4.4 A RNN Model of Planning
Planning with HC-MEC Enables Reconstruction of Sensory Experiences Along Planned Paths. We next test whether the HC-MEC loop enables the network to reconstruct sensory experiences along planned trajectories using intermediate GC states. To this end, we connect an untrained planning subnetwork to a pre-trained HC-MEC model ($r_{\text{mask}}=1.0$, see Suppl. 5.2) and fix all projections from SMCs and HPCs to the planner to zero. This ensures the planner uses only GC representations for planning and controls HC-MEC dynamics by producing inputs to speed and direction cells. SMCs and GCs are initialized to their respective ground truth responses at the start location.
@@ 5 Discussion
Decades of theoretical and computational research have sought to explain how and why hippocampal place cells (HPCs) and grid cells (GCs) emerge. These models address the question “What do they encode?” But an equally important question is “Why does the brain encode?” One answer is that animals develop and maintain place and grid representations to support intrinsically motivated navigation, enabling access to resources critical to survival, such as food, water, and shelter. Thus, we take the perspective: given the known phenomenology of GCs and HPCs, how might they be wired together to support intrinsically motivated navigation?
@@
Previous studies show that HPCs are linked to memory and context but have localized spatial representations lacking the relational structure needed for planning. In contrast, the periodic lattice of GCs supports path planning by generalizing across environments, but weak contextual tuning limits direct recall of GC patterns from sensory cues. Rather than viewing GCs and HPCs as parallel spatial representations, we propose that GCs and spatially modulated cells (SMCs), reflecting sensory observations, form parallel, independent localization systems, each encoding different aspects of space. HPCs then link these representations through auto-association, enabling recovery of one when the other fails and providing a higher-level, abstract spatial representation.
Testing these ideas, we built a single-layer RNN that simultaneously models GCs, HPCs, SMCs, and a planning subnetwork. We showed that auto-association in HPCs, linking MEC sensory (SMCs) and spatial (GCs) representations enables recall of GC patterns from contextual cues. With recalled GC patterns, local planning using only GCs can be learned and generalized to longer trajectories. Finally, HPCs can reconstruct sensory experiences along planned paths, obviating the need for direct planning based solely on HPCs or SMCs. Our model is flexible and accommodates existing theories, as follows. Recurrent connections from MEC cause HPC activity to lag one timestep behind MEC inputs; during pattern completion, HPCs implicitly learn to predict the next timestep’s MEC inputs, and thus can account for the successor representation theory of HPCs [36]. This predictive nature is thus consistent with previous theories that predictive coding in HPCs could support planning [37], but we emphasize that using GCs may be more efficient when it is necessary to shortcut unvisited locations. Lastly, while we propose that the HC–MEC coupling enables planning through recall and reconstruction, a parallel idea was proposed in [70], where the same coupling was argued to support increased episodic memory capacity.
@@
Our theory makes several testable predictions. First, sensory cues should reactivate GC activity associated with a target location, even when the animal is stationary or in a different environment. Second, inhibiting MEC activity should impair path planning and disrupt sequential preplay in the HC. Finally, if hippocampal place cells reconstruct sensory experiences during planning, disrupting MEC-to-HC projections should impair goal-directed navigation, while disrupting HC-to-MEC feedback should reduce planning accuracy by preventing animals from validating planned trajectories internally.

### v2

@@ 3 Recalling MEC Representations from Sensory Observations
We test our first prediction that associating sensory observations (SMC responses) with GC patterns enables recall of those patterns from partial sensory cues. Auto-association arises when input patterns are incomplete or degraded. To model this, we trained nine identical networks differing only in masking ratio $r_{\text{mask}}$ (0.1 to 0.9), which specifies the maximum fraction of head direction, speed, GC, and SMC inputs and initial hidden states randomly set to zero during training. Each model was trained on randomly sampled short trajectories using a fixed $r_{\text{mask}}$, with new masks generated for every trajectory and varying across time and cells. Masks were applied to both inputs and initial states of GCs, SMCs, speed cells, and head direction cells (Suppl. 5.2).
@@
Figure 2a shows the L2 distance between decoded and ground-truth positions over time. The network was queried for 5 seconds (100 timesteps), and all models successfully recalled the goal location with high accuracy. HPCs were identified as neurons with mean firing rates above 0.01 Hz and spatial information content (SIC) above 20 (Suppl. 2). The number of HPCs increased with $r_{\text{mask}}$, and location decoding using HPCs was performed only for models with more than 10 HPCs. We observe a trade-off in which higher $r_{\text{mask}}$ leads to more HPCs and improved decoding accuracy but reduces the network’s ability to recall sensory observations (Figure 2a,b).
@@ 4 Planning with Recalled Representations
The recall experiment shows that auto-associating spatially modulated cells (SMCs) and grid cells (GCs) through hippocampal place cells (HPCs) enables recovery of all cells’ representations from partial sensory cues. Although planning with HPCs and SMCs is feasible [41, 43, 42, 40], their context dependence limits generalization across environments. In contrast, GC patterns encode the geometric structure of space and provide a context-independent substrate for planning. Planning on the grid manifold allows HPC auto-association to reconstruct corresponding SMC representations from intermediate GC states, removing the need to plan directly with HPCs or SMCs. We therefore propose that planning occurs on the grid cell manifold, with sensory details later recovered through hippocampus auto-associations.
@@ 4.1 Decoding Displacement from Grid Cells
Consider a 1D case with $\phi_{c}$ and $\phi_{t}$ as the phase variables of the current and target locations in some module. The phase difference is $\Delta\phi=\phi_{t}-\phi_{c}$, and since $\phi_{c},\phi_{t}\in[0,2\pi)$, we have $\Delta\phi\in(-2\pi,2\pi)$. However, for vector-based navigation, we instead need $\Delta\phi^{*}$ such that $\phi_{t}=\left[\phi_{c}+\Delta\phi^{*}\right]_{2\pi}$, where $\cdot]_{2\pi}$ is an element-wise modulo operation so that $\phi_{t}$ is defined on $0,2\pi)$. Simply using $\Delta\phi$ directly is not sufficient because multiple wrapped phase differences correspond to the same phase $\phi_{t}$, but different physical positions on the torus. Therefore, we restrict $\Delta\phi$ to be defined on $(-\pi,\pi)$ such that the planning mechanism always selects the shortest path on the torus that points to the target phase. The decoded displacement in physical space is then $\hat{d}\in[-\ell/2,\ell/2]$.
@@
Figure 3: (a) The population response of a single grid module forms an $n$-dimensional torus, where multiple phase differences can connect the current and target phases. (b) Unwrapping phases into physical space yields $2^{n}$ candidate displacements; only $\phi_{t}^{*}$ lies within the principal domain $(-\pi,\pi)^{n}$. (c) A Markovian process asymptotically identifies the most likely next phase that moves closer to the target. (d) Renormalizing after each update produces a smooth trajectory from start to target. (e) Illustration of this process in 2D space. (f) Learned connectivity matrix of the planning RNN using only grid cells. (g) Planned trajectories for targets reachable within $10$ and 2020 seconds. Blue and red crosses mark start and target locations; the reference line shows the full trajectory for visualization. The dots represented the decoded locations. (h) A planning network connected to the full HC-MEC, receiving input only from GC and controlling speed and direction, drives SMC responses to update alongside GC, tracing a trajectory closely aligned with the planned GC path.
@@ 4.2 Sequential Planning
By repeatedly applying this update, the influence of the current and target phase spreads through phase space, eventually settling on an intermediate phase that connects start and target (Figure 3c). If the animal selects the phase with the highest value after each update and renormalizes the vector, this process traces a smooth trajectory toward the target (Figure 3d). This approach can be generalized to 2D phase spaces (Figure 3e). In essence, we propose that the animal can decompose long-range planning into a sequence of local steps by encoding a transition probability over phases in a matrix. A readout mechanism can then map these phase transitions into corresponding speeds and directions and subsequently update GC activity toward the target.
@@ 4.3 Combining Decoded Displacement from Multiple Scales
Given the ongoing debate about whether grid cells contribute to long-range planning [64, 65], we focus on mid-range planning, where distances are of the same order of magnitude as the environment size. Suppose two locations in 1D space are separated by a ground truth displacement $d\in\mathbb{R}_{+}$, bounded by half the largest scale ($\ell_{m}/2$). We can always find an index $k$ where $\ell_{1}/2,\dots,\ell_{k}/2\leq d<\ell_{k+1}/2<\dots,\ell_{m}/2$. Given $k$, we call scales $\ell_{k+1},\dots,\ell_{m}$ decodable and scales $\ell_{1},\dots,\ell_{k}$ undercovered. For undercovered scales, phase differences $\Delta\phi$ are wrapped around the torus at least one period of $(-\pi,\pi)$ and may point to the wrong direction. We thus denote the phase difference due to the undercovered scales by $Z_{i}$. If we predict displacement by simply averaging the decoded displacements from all grid scales, the predicted displacement
$\hat{d}=\frac{\ell_{0}}{2\pi m}\cdot\left(\sum_{i=1}^{k}s^{i}\cdot Z_{i}+\sum_{i=k+1}^{m}s^{i}\cdot\Delta\phi_{i}\right).$
In 1D, the remaining distance after taking the predicted displacement is $d_{\text{next}}=d_{\text{current}}-\hat{d}$. For the predicted displacement to always move the animal closer to the target, meaning $d_{\text{next}}<d_{\text{current}}$, it suffices that $m>k+\frac{1-s^{-k}}{s-1}$ (see Suppl. 1). This condition is trivially satisfied in 1D for $s=e$, as $\frac{1-s^{-k}}{s-1}<1$ requiring only $m>k$. In 2D, where the optimal scaling factor $s=\sqrt{e}$, the condition tightens slightly to $m>k+1$. Importantly, as the animal moves closer to the target, more scales become decodable, enabling increasingly accurate predictions that eventually lead to the target. In 2D, planning can be decomposed into two independent processes along non-collinear directions. Although prediction errors in 2D may lead to a suboptimal path, this deviation can be reduced by increasing the number of scales or taking smaller steps along the decoded direction, allowing the animal to gradually approach the target with higher prediction accuracy.
@@ 4.4 A RNN Model of Planning
Figure 4: (a) (Pre-)training process of the BtnkMAE that uses a modification of the Masked AutoEncoder in [48]. Image features are ignored, and only the CLS token is passed to the decoder. The decoder cross-attends to the CLS token with learnable embeddings to reconstruct the image tokens such that each image can be faithfully encoded into 1024 dimensional features. See Suppl. 7. (b) Collection of SMC rate maps from Habitat Sim. (c) Decoding intermediate SMC states during path planning yields images that closely match the expected views along the planned route.
Training the Planning Network Together with HC-MEC Enables Reconstruction of Sensory Experiences Along Planned Paths. We next test whether the HC-MEC loop enables the network to reconstruct sensory experiences along planned trajectories using intermediate GC states. To this end, we train the planning subnetwork together with the entire HC-MEC model that was pre-trained to encode the room during random traversal ($r_{\text{mask}}=1.0$, see Suppl. 5.2). We fix all projections from SMCs and HPCs to the planner to zero. This ensures that the planner uses only GC representations for planning and controls HC-MEC dynamics by producing inputs to speed and direction cells. SMCs and GCs are initialized to their respective ground truth responses at the start location.
@@
Testing REMI for Realistic Navigation Task in Habitat Sim. Finally, we tested whether the proposed framework generalizes to realistic navigation tasks and whether the reconstructed sensory embeddings not only correspond to the correct locations along the planned path but also retain sufficient information to recover the original sensory observations. To this end, we replaced the simulated SMC rate maps $R_{\text{smc}}\in\mathbb{R}^{L_{x}\times L_{y}\times N}$ with visual feature embeddings extracted from the Habitat Synthetic Scenes Dataset (HSSD) [66] in the photorealistic simulator Habitat-Sim [45, 46, 47]. The new SMC rate map collected in Habitat-Sim is $R_{\text{hab}}\in\mathbb{R}^{L_{x}\times L_{y}\times 1024}_{+}$.
To construct $R_{\text{hab}}$, we captured panoramic images $I$ at each spatial location with a fixed orientation and encoded them using a vision encoder that produced 1024-dimensional feature vectors $E(I)\in\mathbb{R}^{1024}_{+}$. We additionally needed a decoder $D(E(I))\approx I$ to transform the encoded image back to the pixel space. For this purpose, we modified a pretrained Masked Autoencoder (MAE) [48] with a ViT-Large backbone, referred to as BtnkMAE. The original MAE produced image features of size $\mathbb{R}^{p\times d}$, where $p$ is the number of patches and $d$ the embedding dimension, which is unsuitable for our HC-MEC model that requires a single vector representation per image. To address this, we pretrained BtnkMAE on ImageNet-1k, retaining only a retrained CLS token after the encoder to represent each image before decoding. The decoder then employed DETR-style cross-attention [67] to reconstruct the image from this single embedding vector (Figure 4a; see also Suppl. 7 for details).
We then repeated the experiment from the previous section with an additional layer normalization to stabilize training. During planning, the HC-MEC updated the SMC region to intermediate states $z_{0},\ldots,z_{T}$, where $z_{i}\in\mathbb{R}^{1024}_{+}$ denotes the population firing statistics of the SMC subregion. Decoding them with $D$ yielded images $I_{0},\ldots,I_{T}$ that closely resembled the expected views at the corresponding planned locations (Figure 4c).
@@ 5 Discussion
Decades of theoretical and computational studies have explored how hippocampal place cells (HPCs) and grid cells (GCs) arise, addressing the question “What do they encode?” An equally important question is “Why does the brain encode?” We suggest that place and grid representations evolved to support internally driven navigation toward vital resources such as food, water, and shelter. Thus, we ask: given their known phenomenology, how might HPCs and GCs be wired to support internally driven navigation?
@@
Previous studies show that HPCs are linked to memory and context but exhibit localized spatial maps that lack the relational structure required for planning, while the periodic lattice of GCs supports path planning and generalization yet exhibits weak contextual tuning that limits direct recall from sensory cues. We propose that GCs and spatially modulated cells (SMCs) form parallel localization systems encoding complementary aspects of space: GCs provide metric structure and SMCs reflect sensory observations. HPCs link these systems through auto-association to enable bidirectional pattern completion. We built a single-layer RNN containing GCs, HPCs, SMCs, and a planning subnetwork to test whether this coupling enables three key capabilities: (1) recall of GC patterns from contextual cues, (2) generalizable GC-based planning to novel routes, and (3) reconstruction of sensory experiences along planned paths, eliminating the need for direct planning on HPC or SMC manifolds.
Our model is flexible and accommodates existing theories. In REMI, HPCs must predict both GC and SMC responses at the next timestep to pattern-complete. Since HPCs are direction-agnostic, accurately reconstructing GC and SMC activity requires encoding possible state transitions. Because adjacent locations tend to have higher transition probabilities [39], our model implicitly reflects the SR framework of HPCs, in which the transition structure is embedded in the pattern-completion dynamics. This predictive nature aligns with previous theories that predictive coding in HPCs supports planning [40], but we emphasize that using GCs may be more efficient when shortcutting unvisited locations. Lastly, while we propose that HC-MEC coupling enables planning through recall and reconstruction, a parallel idea was proposed in [68], where a similar coupling was argued to support greater episodic memory capacity.
@@
Our theory makes several testable predictions. First, sensory cues should reactivate GC activity associated with a target location, even when the animal is stationary or in a different environment. Second, during planning, HC-MEC coupling will induce “look forward” sweeps in MEC spatial representation, resembling recent experimental results [49]. Finally, if hippocampal place cells reconstruct sensory experiences during planning, disrupting MEC-to-HC projections should impair goal-directed navigation, while disrupting HC-to-MEC feedback should reduce planning accuracy by preventing animals from validating planned trajectories internally.